OpenStudy (anonymous):

3y + 2x = 17 -x - 4y = -21 elimination method someone help

OpenStudy (anonymous):

If we multiply the 2nd by 2 and add it to the first we get this: 3y-8y=17-42 so then, solving for y: -5y=-25 y=5 Now we go back to the first equaiton and we get this: 3*5+2x=17 2x=17-15 x=1 so the answer is: x=1, y=5

OpenStudy (anonymous):

thanks you so much

OpenStudy (anonymous):

You are welcome! ;)